Piggly Wiggly has 5,000 shares of 7%, $25 par value, cumulative preferred stock and 150,000 shares of $1 par value common stock outstanding at December 31, 2019 and December 31, 2020. In 2019, a $4,000 dividend was declared and paid. In 2020, $33,000 of dividends are declared and paid. What are the dividends received by the preferred stockholders in 2020 (assuming no dividends in arrears prior to 2019)?

$33,000

$8,750

$13,500

$17,500
Answer #1

(a)Yearly Preferred dividends = number of shares * par value per share * rate*Years

= 5000*25*7%*2= $17500

(b) dividend declared and paid = $4000

Total preferred dividend to be issued (a-b)= $17500 +$4000 = $13500